A uniform electric field is directed to the right and has a magnitude of 8 N/C. Find the magnitude and direction of the force experienced by a 10 nC test charge placed in this field.


1
Expert's answer
2022-03-02T10:14:13-0500
"F=qE=10*10^{-9}*8=80*10^{-9}\:\rm N"

Force is directed to the right.